# AltFS: Agency-light Feature Selection with Large Language Models in Deep Recommender Systems

### v1

@@ 2.5 Optimization
$\text{min}\ \mathcal{L}=-(\boldsymbol{y}_{i}\text{log}(\hat{\boldsymbol{y}}_{i})+(1-\boldsymbol{y}_{i})\text{log}(1-\hat{\boldsymbol{y}}_{i}))$ (9)
@@
2.6 Retraining
Given the well-trained bridge vector $\boldsymbol{w}^{*}$ and weighting gate $\boldsymbol{\hat{w}}^{*}$, we can obtain the final feature importance scores for each feature, which simultaneously consider general world knowledge and recommendation task information.
$\boldsymbol{fi}_{t}^{k}=1-\frac{t}{N}$ (10)
$\boldsymbol{h}_{n}^{*}=\sum\nolimits_{k=0}^{K}\sum\nolimits_{t=0}^{N}\delta(\boldsymbol{F}_{n},f_{t}^{k})\cdot(\boldsymbol{fi}_{t}^{k}\cdot\boldsymbol{\hat{w}}_{k}^{*})$ (11)
where $\boldsymbol{fi}_{t}^{k}$ denotes the feature importance score for feature $\boldsymbol{f}_{t}^{k}$ with $k$th LLM, the score decreases linearly according to the feature’s ranking in the sequence. $\boldsymbol{h}_{n}^{*}$ and $\boldsymbol{\hat{w}}_{k}^{*}$ are the final integrated feature importance score for $n$th feature field and well-trained weighting gate for $k$th LLM expert.
The top-$n$ feature fields with the highest $\boldsymbol{h}_{n}^{*}$ will be selected to retrain DRS from scratch.
Table 1: Comparison results between AltFS and other state-of-the-art baselines on three public datasets. The metrics of the best-performed methods and sub-optimal methods are highlighted in bold fonts and underlined. “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better. $\downarrow$: lower is better.
@@ 2.6 Retraining
AltFS 0.80480 0.52703 0.62015 0.16139 0.78002 0.55801
3 Experiments
3.1 Experimental Settings
3.1.1 Dataset.
@@ 3.1 Experimental Settings
We conduct experiments on three public datasets: Movielens-1M [^2] https://grouplens.org/datasets/movielens/1m/, Aliccp Ma et al. (2018), and Kuairand Gao et al. (2022). These datasets have varying numbers of features (9 in Movielelns-1M, 23 in Aliccp, and 102 in Kuairand). They are chosen to illustrate the effectiveness of our method under different application conditions. Detailed descriptions of datasets are given in Appendix A.2.
3.1.2 Baselines.
We perform experiments with the following baseline methods to demonstrate the advanced performance of our method. (1) Shallow feature selection methods: Lasso Tibshirani (1996), GBDT Friedman (2001), RandomForest Breiman (2001), XGBoost Chen and Guestrin (2016). (2) Gate-based feature selection methods: AutoField Wang et al. (2022), AdaFS Lin et al. (2022), OptFS Lyu et al. (2023), LPFS Guo et al. (2022). (3) Sensitivity-based feature selection methods: Permutation Fisher et al. (2019), SFS Wang et al. (2023), SHARK Zhang et al. (2023). The detailed information of these baselines is introduced in Appendix A.3.
3.1.3 Evaluation Metrics.
@@
To comprehensively evaluate our method, we use the AUC and Logloss metrics in CTR prediction following previous work Wang et al. (2022); Lin et al. (2022). It is worth noting that an increase of 0.001 in AUC or a decrease in Logloss is already a significant improvement in CTR prediction task Song et al. (2019); Wang et al. (2021).
3.1.4 Implementation Details.
@@
We use Adam optimizer with learning rate 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$ for all experiments. The batch size is 4096, the embedding size is 8, the number of LLMs $K$ is 3, the maximum mask ratio is 0.2, and the temperature coefficient $\tau$ is 4.0. To minimize the impact of randomness, we conduct each trial three times and use the average as the final result. The detailed settings are introduced in Appendix A.4.
3.2 Overall Performance
@@
In this section, we compare AltFS with other state-of-the-art baselines on three public datasets. The results are listed in Table 1, and we can derive the following findings: 1) AltFS achieves the best performance across all metrics on three datasets. This superior performance is attributed to AltFS’s unique approach, which does not rely solely on the information provided by agency models to determine feature importance. Instead, it combines both world knowledge and task-specific information to comprehensively assess feature importance. 2) Overall, Gate-based and Sensitivity-based feature selection methods outperform Shallow feature selection methods. This is because Shallow feature selection methods rely on the original classifier, whose performance is generally weaker compared to the deep neural networks employed by Gate-based and Sensitivity-based feature selection methods. 3) The performance improvement of AltFS compared to no selection varies significantly across different datasets. Specifically, the improvement is most pronounced on the Movielens-1M dataset, while it is relatively modest on the Kuairand dataset. This discrepancy is due to the substantial differences in feature quantities among the datasets. The Movielens-1M dataset has fewer features, so removing a certain amount of irrelevant features has a more significant impact and yields greater gains. A case study is given in Appendix A.6.
3.3 Data-scarcity Performance
@@
Table 2: Data-scarcity performance of AltFS and baselines on three datasets. The metrics of best-performed methods are highlighted with bold fonts, and the sub-optimal methods are emphasized with underlined text.
@@
AltFS 0.72372 0.62563 0.55878 0.16735 0.71463 0.66995
In this section, we discuss the performance of AltFS under data-scarcity conditions. To simulate a data-limited scenario, we reduce the training and validation sets of public datasets to 5% of their original size. Table 2 presents the results of AltFS and the baselines across three datasets. Based on Table 2, we can draw the following conclusions: 1) AltFS maintains its stable superiority even in data-scarce scenarios, almost outperforming all baselines across all metrics. This stability is due to AltFS’s ability to optimize the bridge vector with a minimal amount of data, thereby enhancing its effectiveness under limited data conditions. 2) Compared to the results in Table 1, the gap between AltFS and the No Selection approach becomes more pronounced. This indicates that feature selection has a more significant impact on the final results when data is scarce. 3) In the MovieLens-1M dataset, six baselines perform even worse than the No Selection approach. This is likely because, under data-scarce conditions, models are prone to overfitting, which prevents the agency model from providing reliable information for assessing feature importance.
3.4 Transferability Study
Table 3: Transferability study.
Model AUC $\uparrow$ Logloss $\downarrow$
No Selection AltFS No Selection AltFS
FM 0.60667 0.60725 0.16284 0.16283
DeepFM 0.61716 0.61757 0.16203 0.16180
WideDeep 0.62084 0.62212 0.16143 0.16120
DCN 0.62271 0.62322 0.16173 0.16124
In this section, we investigate the transferability of features selected by AltFS. Specifically, whether the selected features remain effective when applied to other deep recommendation backbone models. We apply the features selected by AltFS to four popular deep recommendation models: FM Rendle (2010), DeepFM Guo et al. (2017), Wide&Deep Cheng et al. (2016), and DCN Wang et al. (2017) on the Aliccp dataset. In Table 3, “No Selection” and “AltFS” represent the performance of the model with all available features and with the features selected by AltFS. We can find that: 1) All deep recommendation backbone models achieve improved prediction performance after applying the features selected by AltFS. This demonstrates that the features selected by AltFS possess strong transferability capabilities and can adapt to various deep recommender systems. 2) By comparing the performance improvements across different models after applying the AltFS-selected features, we observe that more complex models (Wide&Deep and DCN) show greater enhancements compared to FM-based models. The possible reason is that more powerful models tend to amplify the impact of the features.
3.5 Ablation Study
In this section, we compare the performance of the following variants on Aliccp to verify the effectiveness of each component: 1) w/o MM: AltFS without multiple LLMs. This variant only considers the feature sequence generated by one advanced LLM (GPT4). 2) w/o IP: AltFS without iterative prompt. This variant prompts LLM to generate feature sequences in one shot, not in an iterative manner. 3) w/o BV: AltFS without bridge vector. This variant directly derives the final feature rankings by averaging the feature sequences generated by LLMs.
Figure 4: Ablation Study.
@@ 3.5 Ablation Study
From Figure 4, we can observe the following conclusions: 1) AltFS outperforms w/o MM. It is because the feature importance ranking from a single advanced LLM has an inherent bias, while AltFS can combine rankings from multiple LLMs to achieve a more robust and effective ranking. 2) AltFS is better than w/o IP. The possible reason is that the iterative prompt method for obtaining feature importance rankings allows the LLM to focus on a more straightforward task each time, resulting in better performance than the one-shot approach. 3) AltFS is superior to w/o BV, indicating the effectiveness of the bridge vector, which can effectively evaluate the correctness of different feature rankings and integrate them. 4) All variants outperform No Selection in all metrics, demonstrating the stability of AltFS.
3.6 Hyperparameter Analysis
@@
Figure 5: Hyperparameter analysis on number of LLMs ($K$) and maximum masking ratio ($\beta$).
@@
In this section, we will investigate the effect of hyperparameters on the performance of the Aliccp dataset. There are two important hyperparameters in AltFS: the number of LLMs ($K$) and the maximum masking ratio ($\beta$). Figure 5 shows the results of AUC on different hyperparameters. The full results containing both AUC and Logloss can be found in Appendix A.5.
3.6.1 Number of LLMs.
@@
As shown in Figure 5(a), we experiment with the number of LLMs from 1 to 4. We prioritize using models with larger parameter sizes, starting with the most powerful model (GPT4). The results lead to the following conclusions: (1) A single strong LLM can provide valuable information for feature selection, with just one LLM outperforming the No selection baseline. (2) The optimal performance is achieved when the number of LLMs is set to 3. This is because the bridge vector can effectively assign weights to the existing feature importance sequences generated by different LLMs according to the current recommendation task. This results in a comprehensive feature importance ranking that integrates both semantic and task-specific aspects. (3) When the number of LLMs exceeds 3, the performance of the method stabilizes, indicating the generalization and stability of AltFS, with 3 being the most cost-effective hyperparameter choice.
3.6.2 Maximum Masking Ratio.
@@
As illustrated in Figure 5(b), we set the maximum masking ratio from 0 to 1 and freeze other hyperparameters. We can draw the following conclusions: 1) The overall AUC trend exhibits an initial increase followed by a decrease. This is because when $\beta$ approaches 0, no features will be masked, resulting in the final feature ranking being a simple average of the sequences generated by LLMs. Conversely, when $\beta$ approaches 1, the most important features in the sequences generated by the LLMs tend to be similar at the top, while the differences among the features in the lower ranks are gradually diminished as beta increases. This weakens the effectiveness of feature selection. 2) We can also observe that the feature selection performance remains consistently exceptional when $\beta$ is between 0.2 and 0.5, highlighting the robustness and generalization capability of AltFS.
4 Related Work

### v2

@@ 2.5. Optimization
(9) $\text{min}\ \mathcal{L}=-(\boldsymbol{y}_{i}\text{log}(\hat{\boldsymbol{y}}_{i})+(1-\boldsymbol{y}_{i})\text{log}(1-\hat{\boldsymbol{y}}_{i}))$
@@
2.6. Retraining
In this subsection, we describe how to select informative feature fields based on the learned fusion weights and retrain the deep recommender system.
Given the well-trained bridge vector $\boldsymbol{w}^{*}$, we compute the final fusion weights $\hat{\boldsymbol{w}}^{*}=\text{Softmax}(\boldsymbol{w}^{*}\cdot\exp(\tau))$, which reflect the reliability of each LLM expert. For each feature ranked at position $t$ by the $k$-th LLM, we assign a linearly decaying importance score:
(10) $\text{fi}_{t}^{k}=1-\frac{t}{N}$
where $\text{fi}_{t}^{k}$ denotes the relative importance of the $t$-th ranked feature in the $k$-th LLM ranking. We then aggregate the importance scores across all experts to obtain the final importance score $\boldsymbol{h}_{n}^{*}$ for the $n$-th feature field:
(11) $\boldsymbol{h}_{n}^{*}=\sum_{k=0}^{K}\hat{w}_{k}^{*}\cdot\sum_{t=0}^{N}\mathbb{I}[f_{t}^{k}=\boldsymbol{F}_{n}]\cdot\text{fi}_{t}^{k}$
where $\hat{w}_{k}^{*}$ is the learned fusion weight for the $k$-th LLM. $\mathbb{I}[f_{t}^{k}=\boldsymbol{F}_{n}]$ is the indicator function that equals 1 if the $n$-th feature field appears at position $t$ in the $k$-th ranking. $\boldsymbol{h}_{n}^{*}$ is the final integrated importance score for feature field $\boldsymbol{F}_{n}$.
We then select the top-$d$ feature fields with the highest $\boldsymbol{h}_{n}^{*}$ values and retrain the deep recommender system from scratch using only the selected features.
3. Experiments
To achieve a comprehensive understanding of SELF, we aim to answer the following research questions:
  • RQ1: How does the performance of SELF compare to the other state-of-the-art feature selection methods in DRS?
  • RQ2: How does the transferability of SELF to other backbone deep recommendation models?
  • RQ3: What’s the specific effect of the proposed components?
  • RQ4: How SELF performs under data-scarcity scenarios?
  • RQ5: How do hyperparameters influence performance?
  • RQ6: How SELF performs in real industrial environments?
3.1. Experimental Settings
Table 1. Dataset statistics
Dataset Movielens-1M Aliccp Kuairand
Interactions 1,000,209 85,316,519 1,436,609
Users 6,040 238,635 27,285
Items 3,706 467,298 7,551
Interaction Type Rating (1-5) Click Click
Feature Num 9 23 96
Table 2. Comparison results between SELF and other state-of-the-art baselines on three public datasets. The metrics of the best-performed methods and sub-optimal methods are highlighted in bold fonts and underlined. $\uparrow$ denotes the higher is better. $\downarrow$ denotes the lower is better.
@@ 3.1. Experimental Settings
[…]
3.1.1. Dataset.
@@
We conduct experiments on three public datasets: Movielens-1M [^1] https://grouplens.org/datasets/movielens/1m/, Aliccp (Ma et al., 2018), and Kuairand (Gao et al., 2022). These datasets have varying numbers of features (9 in Movielens-1M, 23 in Aliccp, and 96 in Kuairand). They are chosen to illustrate the effectiveness of our method under different application conditions. The statistics of datasets are illustrated in Table 1 and the detailed descriptions of the datasets are given in Appendix A.1.
3.1.2. Baselines.
We perform experiments with the following baselines to demonstrate the advanced performance of SELF.
Shallow Feature Selection Methods:
  • Lasso (Tibshirani, 1996). Lasso is the Least Absolute Shrinkage and Selection Operator, which effectively combines variable selection and regularization to enhance model performance.
  • GBDT (Friedman, 2001). Gradient Boosted Decision Trees (GBDT) is a highly effective machine learning algorithm that leverages the strengths of decision trees in combination with the boosting technique. It constructs an ensemble of decision trees, with each tree iteratively correcting the errors of its predecessors.
  • Random Forest (Breiman, 2001). Random Forest determines feature importance by measuring the extent to which each feature reduces impurity within a decision tree.
  • XGBoost (Chen and Guestrin, 2016). XGBoost is a highly efficient and scalable machine learning algorithm built on the principles of gradient boosting. XGBoost determines feature importance by assessing the impact of each feature on the model’s final predictive performance.
Gate-based Feature Selection Methods:
  • AutoField (Wang et al., 2022). AutoField is the first work in the recommender systems domain to focus on feature selection within DRS (Deep Recommender Systems). It introduces an advanced controller network that trains two parameters for each feature field, representing the likelihood of selecting that feature.
  • AdaFS (Lin et al., 2022). AdaFS refines feature selection further along the lines of AutoField. Recognizing that feature importance varies significantly across different samples, AdaFS focuses on generating dynamic gating weights for each sample, enabling the model to adaptively select features.
  • OptFS (Lyu et al., 2023). OptFS differs from previous work by focusing on feature value selection. Through the learning of a gating network, OptFS can significantly reduce the number of feature values while simultaneously enhancing the performance.
  • LPFS (Guo et al., 2022). LPFS critiques previous approaches that determine feature importance based on the magnitude of gating values. It introduces a smoothed-$l^{0}$ function that effectively selects features in a single stage.
Sensitivity-based Feature Selection Methods:
  • Permutation (Fisher et al., 2019). This method assesses the importance of a feature by disrupting the values within a specific feature field and observing whether this permutation significantly impacts the model’s predictive performance.
  • SFS (Wang et al., 2023). SFS assigns feature importance weights by examining the gradient of the gating vector on a small subset of data.
  • SHARK (Zhang et al., 2023). SHARK uses the first-order component of the Taylor expansion as a measure of feature importance. It improves model efficiency and performance by removing less important features from the embedding table.
3.1.3. Evaluation Metrics.
@@
To comprehensively evaluate SELF, we use the AUC and Logloss metrics in CTR prediction following previous work (Wang et al., 2022; Lin et al., 2022). It is worth noting that an increase of 0.001 in AUC or a decrease in Logloss is already a significant improvement in the CTR prediction task (Song et al., 2019; Wang et al., 2021).
3.1.4. Implementation Details.
@@
We use Adam (Kingma, 2014) optimizer with learning rate 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$ for all experiments. We set the batch size to 4096 and the embedding dimension to 8. The number of LLMs is $K=3$ (GPT4, GPT4o, and GPT-3.5), the maximum masking ratio is 0.2, and the temperature coefficient is $\tau=4.0$. To mitigate the effect of randomness, each experiment is repeated three times, and the average performance is reported. Detailed implementation settings are provided in Appendix A.2.
3.2. Overall Performance (RQ1)
@@
In this section, we compare SELF with other state-of-the-art baselines on three public datasets. The results are listed in Table 2, and we can derive the following findings: 1) SELF achieves the best performance across all metrics on three datasets. This superior performance is attributed to SELF’s unique approach, which does not rely solely on the information provided by surrogate models to determine feature importance. Instead, it combines both world knowledge and task-specific information to comprehensively assess feature importance. 2) Overall, Gate-based and Sensitivity-based feature selection methods outperform Shallow feature selection methods. This is because Shallow feature selection methods rely on the original classifier, whose performance is generally weaker compared to the deep neural networks employed by Gate-based and Sensitivity-based feature selection methods. 3) The performance improvement of SELF compared to no selection varies significantly across different datasets. Specifically, the improvement is most pronounced on the Movielens-1M dataset, while it is relatively modest on the Kuairand dataset. This discrepancy is due to the substantial differences in feature quantities among the datasets. The Movielens-1M dataset has fewer features, so removing a certain amount of irrelevant features has a more significant impact and yields greater gains.
3.3. Transferability Study (RQ2)
In this section, we investigate the transferability of features selected by SELF. Specifically, whether the selected features remain effective when applied to other deep recommendation backbone models. We apply the features selected by SELF to four popular deep recommendation models: FM (Rendle, 2010), DeepFM (Guo et al., 2017), Wide&Deep (Cheng et al., 2016), and DCN (Wang et al., 2017) on the Aliccp dataset. In Table 3, “No Selection” and “SELF” represent the performance of the model with all available features and with the features selected by SELF. We can find that: 1) All deep recommendation backbone models achieve improved prediction performance after applying the features selected by SELF. This demonstrates that the features selected by SELF possess strong transferability capabilities and can adapt to various deep recommender systems. 2) By comparing the performance improvements across different models after applying the SELF-selected features, we observe that more complex models (Wide&Deep and DCN) show greater enhancements compared to FM-based models. The possible reason is that more powerful models tend to amplify the impact of feature selection.
Table 3. Transferability study.
Model AUC $\uparrow$ Logloss $\downarrow$
No Selection SELF No Selection SELF
FM 0.60667 0.60725 0.16284 0.16283
DeepFM 0.61716 0.61757 0.16203 0.16180
WideDeep 0.62084 0.62212 0.16143 0.16120
DCN 0.62271 0.62322 0.16173 0.16124
3.4. Ablation Study (RQ3)
Figure 3. Ablation Study.
@@ 3.4. Ablation Study (RQ3)
Table 4. Data-scarcity performance of SELF and baselines on three datasets. The metrics of best-performed methods are highlighted with bold fonts, and the sub-optimal methods are emphasized with underlined text.
@@
[…]
@@
In this section, we compare the performance of the following variants on Aliccp to verify the effectiveness of each component: 1) w/o MM: SELF without multiple LLMs. This variant only considers the feature sequence generated by one advanced LLM (GPT4222https://openai.com/index/gpt-4/). 2) w/o IP: SELF without iterative prompt. This variant prompts the LLM to generate feature sequences in one shot, not in an iterative manner. 3) w/o BV: SELF without bridge vector. This variant directly derives the final feature rankings by averaging the feature sequences generated by LLMs. From Figure 3, we can observe the following conclusions: 1) SELF outperforms w/o MM. It is because the feature importance ranking from a single advanced LLM has an inherent bias, while SELF can combine rankings from multiple LLMs to achieve a more robust and effective ranking. 2) SELF is better than w/o IP. The possible reason is that the iterative prompt method for obtaining feature importance rankings allows the LLM to focus on a more straightforward task each time, resulting in better performance than the one-shot approach. 3) SELF is superior to w/o BV, indicating the effectiveness of the bridge vector, which can effectively evaluate the correctness of different feature rankings and integrate them. 4) All variants outperform “No Selection” in all metrics, demonstrating the stability of SELF.
3.5. Data-scarcity Performance (RQ4)
In this section, we discuss the performance of SELF under data-scarcity conditions. To simulate a data-limited scenario, we reduce the training and validation sets of public datasets to 5% of their original size. Table 4 presents the results of SELF and the baselines across three datasets. Based on Table 4, we can draw the following conclusions: 1) SELF maintains its stable superiority even in data-scarce scenarios, almost outperforming all baselines across all metrics. This stability is due to SELF’s ability to optimize the bridge vector with a minimal amount of data, thereby enhancing its effectiveness under limited data conditions. 2) Compared to the results in Table 2, the gap between SELF and the No Selection approach becomes more pronounced. This indicates that feature selection has a more significant impact on the final results when data is scarce. 3) In the MovieLens-1M dataset, six baselines perform even worse than the No Selection approach. This is likely because, under data-scarce conditions, models are prone to overfitting, which prevents the surrogate model from providing reliable information for assessing feature importance.
3.6. Hyperparameter Analysis (RQ5)
@@
Figure 4. Hyperparameter analysis on number of LLMs ($K$) and maximum masking ratio ($\beta$).
@@
In this section, we will investigate the effect of hyperparameters on the performance of the Aliccp dataset. There are two important hyperparameters in SELF: the number of LLMs ($K$) and the maximum masking ratio ($\beta$). Figure 4 shows the results of AUC and Logloss on different hyperparameters.
3.6.1. Number of LLMs.
@@
As shown in the left portion of Figure 4, we experiment with the number of LLMs from 1 to 4 (GPT4, GPT4o, GPT3.5, and Gemini-1.5-Pro). We prioritize using models with larger parameter sizes, starting with the most powerful model. The results lead to the following conclusions: (1) A single strong LLM can provide valuable information for feature selection, with just one LLM outperforming the No selection baseline. (2) The optimal performance is achieved when the number of LLMs is set to 3. This is because the bridge vector can effectively assign weights to the existing feature importance sequences generated by different LLMs according to the current recommendation task. This results in a comprehensive feature importance ranking that integrates both semantic and task-specific aspects. (3) When the number of LLMs exceeds 3, the performance of the method stabilizes, indicating the generalization and stability of SELF, with 3 being the most cost-effective hyperparameter choice.
3.6.2. Maximum Masking Ratio.
@@
As illustrated in the right portion of Figure 4, we set the maximum masking ratio from 0 to 1 and freeze other hyperparameters. We can draw the following conclusions: 1) The overall AUC trend exhibits an initial increase followed by a decrease. This is because when $\beta$ approaches 0, no features will be masked, resulting in the final feature ranking being a simple average of the sequences generated by LLMs. Conversely, when $\beta$ approaches 1, the most important features in the sequences generated by the LLMs tend to be similar at the top, while the differences among the features in the lower ranks are gradually diminished as beta increases. This weakens the effectiveness of feature selection. 2) We can also observe that the feature selection performance remains consistently exceptional when $\beta$ is between 0.2 and 0.5, highlighting the robustness and generalization capability of SELF.
3.7. Online Experiments (RQ6)
We deployed SELF on a large-scale industrial search advertising platform that serves billions of users daily. Users interact with the platform by submitting search queries and receiving responses about relevant applications and associated advertisements. In our system, ads from different advertisers are divided into two stages based on their onboarding time: a learning phase and a stable phase. Ads in the learning phase typically suffer from limited user feedback due to their short exposure duration, leading to low-quality and sparse training samples. Under such conditions, traditional data-driven feature selection methods—such as AutoField and SFS—struggle to identify high-quality features, as they heavily depend on the availability of abundant, high-quality labeled data.
To address the challenges in the learning phase, we deployed the SELF algorithm to assist with feature selection under limited data conditions. Specifically, we provided the Qwen family models with business-related features and their corresponding descriptions, along with a clear explanation of the recommendation task. This enabled the LLMs to generate initial feature importance rankings based on semantic reasoning. We then refine these rankings using the limited training data available in the learning phase. Based on the resulting importance scores, we removed the 12 least important features and deployed the resulting model as the experimental group. All other parameters were held constant, isolating the impact of feature selection. Compared to the baseline model, the experimental model achieved a 0.63% improvement in RPM (Revenue per Mille), a 3.01% increase in CTR (Click-Through Rate), and a 13.6% reduction in online inference latency. This optimized feature set has since been adopted to serve the entire production traffic.
3.8. Case Study
Figure 5. (a) Cramér’s V matrix and (b) AUC performance on Movielens-1M.
In this section, we present a specific example to illustrate the effectiveness of LLM reasoning in the feature selection process. Figure 5(a) shows the Cramér’s V index of various features in the MovieLens-1M dataset, representing the correlation between different features. The darker the blue, the stronger the correlation between the two features. We observe that “title” and “movie_id” are highly correlated, as indicated by the red box in the figure. Additionally, Figure 5(b) displays the AUC performance on the MovieLens-1M dataset for the best-performing baseline LPFS, LPFS-t (which removes the title feature based on LPFS), and SELF. We find that removing the title feature results in a significant improvement for LPFS, indicating that the strong correlation between “title” and “movie_id” indeed hinders model training. Finally, we examine the feature importance ranking from LPFS: [“movie_id”, “user_id”, “title”, “genres”, “zip”, “gender”, “age”, “occupation”, “timestamp”]. LPFS assigns high importance to both “movie_id” and “title”, likely because the model optimization process treats these two features as roughly equally important, overlooking the issue of information redundancy between them. However, SELF, by leveraging LLM reasoning, effectively identifies this issue. Below is an analysis extracted from the LLM response during prompt iteration regarding the title feature: “While the movie title might have some influence, it’s likely highly correlated with movie_id and thus doesn’t add significant independent information.” Consequently, a more accurate and effective feature ranking is obtained: [“user_id”, “movie_id”, “genres”, “age”, “gender”, “timestamp”, “occupation”, “zip”, “title”]. This demonstrates the effectiveness of SELF in handling feature importance extraction.
4. Related Work
4.1. Feature Selection for DRS.